• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 16:19:25 on July 26, 2006

    Entry number 175389

    keyword=Beam Time Accounting
    Hall A

    07/26/06 0800-1600
    Experiment #: CSR
    Shift Leader: J. P. Chen
    Shift Workers: G. Kumbartzki, R. Sparks

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    0.500.710.050.000.000.000.000.001.26
    Beam Available
    but Not in Use
    0.290.180.951.001.001.001.001.006.43
    Beam Not Available
    or Unacceptable
    0.210.100.000.000.000.000.000.000.31
    Accelerator Con-
    figuration Change
    0.000.000.000.000.000.000.000.000.00
    Experiment
    Ready
    0.720.820.050.000.000.000.000.001.58
    Planned Config-
    uration Change
    0.280.180.951.001.001.001.001.006.42
    Unplanned
    Experiment Down
    0.000.000.000.000.000.000.000.000.00
    Hall not
    ready reasons
    Target0.000.000.000.000.000.000.000.000.00
    Data Acq0.000.000.000.000.000.000.000.000.00
    Other0.000.000.000.000.000.000.000.000.00
    HRSL0.000.000.000.000.000.000.000.000.00
    HRSR0.000.000.000.000.000.000.000.000.00
    HALL_OFF0.000.000.000.000.000.000.000.000.00

    Comments: 8:00 beam on and off. Take data for CRS test. 120 degree, kin. 1, and kin 2. current low (only 5 uA), then raised to 8 uA. Later to 20 uA when Hall C was down. 10:00 CSR test completed. Beam off. Hall in restricted access.

    Data Entry by: jpchen